OpenStudy (anonymous):

solve the Equation for Y 4x-10y=12

OpenStudy (phi):

to solve for y in 4x-10y=12 you do it in steps. One step might be, add +10y to both sides 4x-10y + 10y = 12+10y now simplify the left side. can you do that ?

OpenStudy (nurali):

4x-10y=12 4x-12=10y y=4x-12/10 y=2(2x-6)/10 y=2x-6/5